OpenStudy (anonymous):

Solve the equation in two different ways: x+0.7=1-0.2x.

OpenStudy (anonymous):

\[x+0.7=1-0.2x\] Subtract x from left side, right side. \[0.7=1-0.2x-x\] Simplify. \[0.7=1-1.2x\] subtract 1 to left side, subtract 1 from right side. \[1.7=-1.2x\] \[\frac{ 0.3 }{ -1.2 } = x\] Now just divide.